# Bayesian Experimental Design via Contrastive Diffusions

### v1

@@ 6.2 Sources location finding
${\boldsymbol{\theta}}^{(t+1)}_{i}$ $={\boldsymbol{\theta}}^{(t)}_{i}-\gamma_{t}\nabla_{\boldsymbol{\theta}}\log p({\boldsymbol{\theta}}^{(t)}_{i}|{\mathb{D}}_{k-1})+\sqrt{2\gamma_{t}}{\mathb{B}}_{{\boldsymbol{\theta}},t}\quad\text{and}\quad{\mathb{y}}^{(t+1)}_{i}\sim p({\mathb{y}}|{{\boldsymbol{\theta}}}^{(t+1)}_{i},{\boldsymbol{\xi}}_{t})$ (20)
${{\boldsymbol{\theta}}^{\prime}}^{(t+1)}_{j}$ $={{\boldsymbol{\theta}}^{\prime}}^{(t)}_{j}-\gamma^{\prime}_{t}\sum_{i=1}^{N}\nu_{i}\nabla_{\boldsymbol{\theta}}\log p({{\boldsymbol{\theta}}^{\prime}}^{(t)}_{j}|{\mathb{y}}^{(t+1)}_{i},{\boldsymbol{\xi}}_{t},{\mathb{D}}_{k-1})+\sqrt{2\gamma^{\prime}_{t}}{\mathb{B}}_{{\boldsymbol{\theta}}^{\prime},t}\;.$ (21)
@@
Figure 3: Source location. Median over 100 rollouts for SPCE, L2 Wasserstein distance (logarithmic scale) and SNMC with respect to the number of experiments $k$. Number of samples $N+M=400$
@@ 6.3 Image reconstruction with Diffusion models
We build an artificial experimental design task to illustrate the ability of our method to handle design parameters related to inverse problems where the parameter ${\boldsymbol{\theta}}$ of interest is high dimensional. More specifically, we consider the task that consists of recovering an hidden image from only partial observations of its pixels. The image to be recovered is denoted by ${\boldsymbol{\theta}}$. An experiment corresponds to the choice of a pixel ${\boldsymbol{\xi}}$ around which an observation mask is centered and the image becomes visible. The measured observation ${\mathb{y}}$ is then a masked version of ${\boldsymbol{\theta}}$. The likelihood derives from the model ${\mathb{Y}}={\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}+{\boldsymbol{\eta}}$ where ${\mathb{A}}_{\boldsymbol{\xi}}$ is a square mask centered at ${\boldsymbol{\xi}}$ and ${\boldsymbol{\eta}}$ some Gaussian variable. For the image prior, we consider a diffusion model trained for generation of the MNIST dataset (LeCun et al.,, 1998). The goal is thus to select sequentially the best central pixel locations for $7\times 7$ masks so as to reconstruct an entire $28\times 28$ MNIST image in the smallest number of experiments. The smaller the mask the more interesting it becomes to optimally select the mask centers. Algorithm 2 is used with diffusion-based sampling operators specified in Appendix F.2.2. The gain in optimizing the mask placements is illustrated in Figure 1 and Appendix Figure 8. Progressive image reconstructions are shown in Figure 4 and Appendix Figures 6 and 7. The digit to be recovered is shown in the 1st column. The successively selected masks are shown (red line squares) in the 2nd column with the resulting gradually discovered part of the image. The reconstruction per se can be estimated from the posterior samples shown in the last 16 columns. At each experiment, the upper sub-row shows the 16 most-likely reconstructed images, while the lower sub-row shows the 16 less probable ones. As the number of experiments increases the posterior samples gradually concentrate on the right digit.
@@ 7 Conclusion
We presented a new approach to gradient-based BOED that allows very efficient implementations. Introducing the concept of expected posterior distribution, we proposed a new EIG gradient estimator, which allows the EIG maximization to be performed in a bi-level optimization framework and without resorting to lower bounds. The performance was illustrated in a traditional density-based setting with superior accuracy and lower computational cost compared to state-of-the-art methods. In addition, the possibility of our method to also handle data-based sampling represents, to our knowledge, the first extension of BOED to diffusion-based generative models. By integrating the highly successful framework of diffusion models for our sampling operators, we were able to optimize a design parameter ${\boldsymbol{\xi}}$ concurrently with the diffusion process. This was illustrated in a new application for BOED involving high dimensional image parameters. The foundation of our approach lies on conditional diffusion models and their application to inverse problems. Thanks to this advancement, there are as many new potential applications of BOED as there are trained diffusion models for specific inverse problem tasks. Currently, most models assume a linear forward model. However, the non-linear setting is an active field of research, and advancements in this area could be directly applied to our framework. The applicability of our method could also be extended by considering settings with no explicit expression of the likelihood and investigating simulation-based inference such as developed by Ivanova et al., (2021); Kleinegesse and Gutmann, (2021); Kleinegesse et al., (2020). In addition, although in density-based BOED, we have shown that greedy approaches could outperform long-sighted reinforcement learning procedures, in a data-based setting, it would be interesting to investigate an extension to non myopic approaches such as Iqbal et al., (2024).
@@ Appendix B Logarithmic pooling as a good importance sampling proposal
When considering importance sampling with a proposal distribution $q$ and a target distribution $p$, Chatterjee and Diaconis, (2018) proved that under certain conditions, the number of simulation draws required for both importance sampling and self normalized importance sampling (SNIS) estimators to have small L1 error with high probability was roughly $\exp(\text{KL}(p,q))$, see Theorem 1.2 in Chatterjee and Diaconis, (2018) for SNIS. Similarly, selecting a proposal distribution which minimizes the importance sampling estimator variance is equivalent to finding a distribution with small $\chi^{2}$-distance to $p$, see e.g. Appendix E of Minka, (2005). More generally, finding a good proposal $q$ is linked to the problem of minimizing $\alpha$-divergences or $f$-divergence between $p$ and $q$, which are jointly convex in $p$ and $q$, see Minka, (2005). In this work, we consider $\text{KL}(q,p)$ as a measure of proximity between $p$ and $q$. This choice is ultimately arbitrary but has the advantage of leading to an interpretable proposal with interesting sampling properties. To justify the expected posterior $q_{{\boldsymbol{\xi}},N}$ in (9) and its use in (10), we then use Lemma 2 below to show that for $\sum_{i=1}^{N}\nu_{i}=1$, the distribution $q^{*}$ that minimizes the weighted sum of the KL against each posterior $p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$, i.e. $\sum_{i=1}^{N}\nu_{i}\text{KL}(q,p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}}))$ is
@@
Example: as already mentioned our expected posterior $q_{{\boldsymbol{\xi}},N}$ corresponds to the application of this result to $\rho=\sum_{i=1}^{N}\nu_{i}\delta_{{\mathb{y}}_{i}}$ with $\sum_{i=1}^{N}\nu_{i}=1$.
@@
Remark 3: The expected posterior distribution can also be recovered as a constrained mean field solution. Indeed, it is easy to show that $q^{*}$ is also the measure that minimizes the KL between the joint distribution and a product form approximation where one of the factor is fixed to $\rho({\mathb{y}})$,
@@ C.1 Denoising diffusion models
where ${\boldsymbol{\epsilon}}\sim\mathcal{N}(\mathbf{0},{\mathb{I}})$ is a standard Gaussian random variable. Samples from $p_{0}$ are transformed to samples from a standard Gaussian distribution after some large time $T$.
@@
where $p_{t}$ is the distribution of $\tilde{{\boldsymbol{\theta}}}^{(t)}$ from (34). Solving this reverse SDE, the distribution of ${\boldsymbol{\theta}}^{(T)}$ is closed to $p_{0}$ for large $T$, which allows approximate sampling from $p_{0}$.
The score function $\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)})$ of the noisy data distribution at time $t$ is intractable and is then estimated by learning a neural network $s_{\phi}({\boldsymbol{\theta}},t)$ with parameters $\phi$. Score matching (Hyvärinen,, 2005) is a method to train $s_{\phi}$ by minimizing the following loss:
@@
${\boldsymbol{\theta}}^{(t-\Delta t)}={\boldsymbol{\theta}}^{(t)}+\frac{\beta(t)}{2}{\boldsymbol{\theta}}^{(t)}\Delta t+\beta(t)s_{\phi}({\boldsymbol{\theta}}^{(t)},t)\Delta t+\sqrt{\beta(t)\Delta t}{\boldsymbol{\epsilon}}\;.$ (40)
@@ C.2 Conditional Diffusion Models
$d{\boldsymbol{\theta}}^{(t)}=\left[-\frac{\beta(t)}{2}{\boldsymbol{\theta}}^{(t)}-\beta(t)\nabla_{\boldsymbol{\theta}}\log p({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})\right]dt+\sqrt{\beta(t)}d{\mathb{B}}_{t},$ (42)
with the usual score $\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)})$ replaced by the conditionnal score $\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})$. The main objective of conditional SDE is to generate samples from the conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ without retraining a new neural network $s_{\phi}$ for the new conditional score. By noting that the conditional score can be written as:
$\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})=\nabla_{\boldsymbol{\theta}}\log p_{t}({\mathb{y}}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})+\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)})$ (43)
we can leverage a pre-computed neural network $s_{\phi}$ that was trained to estimate the score $\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}}^{(t)})$ in the unconditional case. If we know how to evaluate the first term $\nabla_{\boldsymbol{\theta}}\log p_{t}({\mathb{y}}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$, we can then run the reverse process (42) to generate samples from the conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$. Unfortunately this term does not have a closed form expression. As a solution, Dou and Song, (2024) propose to approximate the intractable $\nabla_{\boldsymbol{\theta}}\log p_{t}({\mathb{y}}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$ by the tractable $\nabla_{\boldsymbol{\theta}}\log p_{t}({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$ where ${\mathb{y}}^{(t)}$ is a noisy version of ${\mathb{y}}$ at time $t$. Then, the following backward SDE can be run to generate samples from the conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$:
$d{\boldsymbol{\theta}}^{(t)}=\left[-\frac{\beta(t)}{2}{\boldsymbol{\theta}}^{(t)}-\beta(t)\nabla_{\theta}\log p_{t}({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})-\beta(t)\nabla_{\theta}\log p_{t}({\boldsymbol{\theta}}^{(t)})\right]dt+\sqrt{\beta(t)}d{\mathb{B}}_{t}$ (44)
@@
${\mathb{y}}^{(t)}={\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}^{(t)}+\sqrt{\bar{\alpha}_{t}}{\boldsymbol{\eta}}\;.$ (46)
We can then evaluate $\nabla_{\boldsymbol{\theta}}\log p_{t}({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})$ as :
$\nabla_{\boldsymbol{\theta}}\log p_{t}({\mathb{y}}^{(t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})=\frac{1}{\sigma^{2}\bar{\alpha}_{t}}{\mathb{A}}_{\boldsymbol{\xi}}^{T}({\mathb{y}}^{(t)}-{\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}^{(t)})\;.$ (47)

### v2

@@ 6.2 Sources location finding
${\boldsymbol{\theta}}^{(t+1)}_{i}$ $={\boldsymbol{\theta}}^{(t)}_{i}+\gamma_{t}\nabla_{\boldsymbol{\theta}}\log p({\boldsymbol{\theta}}^{(t)}_{i}|{\mathb{D}}_{k-1})+\sqrt{2\gamma_{t}}{\mathb{B}}_{{\boldsymbol{\theta}},t}\quad\text{and}\quad{\mathb{y}}^{(t+1)}_{i}\sim p({\mathb{y}}|{{\boldsymbol{\theta}}}^{(t+1)}_{i},{\boldsymbol{\xi}}_{t})$ (20)
${{\boldsymbol{\theta}}^{\prime}}^{(t+1)}_{j}$ $={{\boldsymbol{\theta}}^{\prime}}^{(t)}_{j}+\gamma^{\prime}_{t}\sum_{i=1}^{N}\nu_{i}\nabla_{\boldsymbol{\theta}}\log p({{\boldsymbol{\theta}}^{\prime}}^{(t)}_{j}|{\mathb{y}}^{(t+1)}_{i},{\boldsymbol{\xi}}_{t},{\mathb{D}}_{k-1})+\sqrt{2\gamma^{\prime}_{t}}{\mathb{B}}_{{\boldsymbol{\theta}}^{\prime},t}\;.$ (21)
@@
Figure 3: Source location. Median and standard error over 100 rollouts for SPCE, L2 Wasserstein distance (log-scale), SNMC with respect to number of experiments $k$. Number of samples N+M=400.
@@ 6.3 Image reconstruction with Diffusion models
We build an artificial experimental design task to illustrate the ability of our method to handle design parameters related to inverse problems with a high dimensional parameter ${\boldsymbol{\theta}}$. We consider the task of recovering an hidden image from only partial observations of its pixels. The image to be recovered is denoted by ${\boldsymbol{\theta}}$. An experiment corresponds to the choice of a pixel ${\boldsymbol{\xi}}$ around which an observation mask is centered and the image becomes visible. The measured observation ${\mathb{y}}$ is then a masked version of ${\boldsymbol{\theta}}$. The likelihood derives from the model ${\mathb{Y}}={\mathb{A}}_{\boldsymbol{\xi}}{\boldsymbol{\theta}}+{\boldsymbol{\eta}}$ where ${\mathb{A}}_{\boldsymbol{\xi}}$ is a square mask centered at ${\boldsymbol{\xi}}$ and ${\boldsymbol{\eta}}$ some Gaussian variable. For the image prior, we consider a diffusion model trained for generation of the MNIST dataset (LeCun et al.,, 1998). The goal is thus to select sequentially the best central pixel locations for $7\times 7$ masks so as to reconstruct an entire $28\times 28$ MNIST image in the smallest number of experiments. The smaller the mask the more interesting it becomes to optimally select the mask centers. Algorithm 2 is used with diffusion-based sampling operators specified in Appendix F.2.2. The gain in optimizing the mask placements is illustrated in Figure 1 and Appendix Figure 9. It is confirmed quantitatively in Table 1, which reports reconstruction quality as measured by the structural similarity index measure (SSIM) (Wang et al.,, 2004), details in Appendix F.2.2. Progressive reconstructions are shown in Figures 4, 7 and 8. The digit to be recovered is shown in the 1st column. The successively selected masks are shown (red line squares) in the 2nd column with the resulting gradually discovered part of the image. The reconstruction per se can be estimated from the posterior samples shown in the last 16 columns. At each experiment, the upper sub-row shows the 16 most-likely reconstructed images, while the lower sub-row shows the 16 less probable ones. As the number of experiments increases the posterior samples gradually concentrate on the right digit.
@@ 7 Conclusion
We presented a new approach, CoDiff, to gradient-based BOED that allows very efficient implementations. The performance was illustrated in a traditional density-based setting with superior accuracy and lower computational cost compared to state-of-the-art methods. In addition, the possibility of our method to also handle data-based sampling represents, to our knowledge, the first extension of BOED to diffusion-based generative models. By integrating the highly successful framework of diffusion models for our sampling operators, we were able to optimize a design parameter ${\boldsymbol{\xi}}$ concurrently with the diffusion process. This was illustrated in a new application for BOED involving high dimensional image parameters. The foundation of our approach lies on a new EIG gradient estimator, bi-level optimization, conditional diffusion models and their application to inverse problems. Thanks to this advancement, there are as many new potential applications of BOED as there are trained diffusion models for specific inverse problem tasks. Current limitations include that CoDiff remains a greedy approach, that it requires an explicit expression of the likelihood and that when using diffusions to address inverse problems only linear forward models are currently handled. However, the non-linear setting is an active field of research, and advancements in this area could be directly applied to our framework. The applicability of our method could also be extended by considering settings with no explicit expression of the likelihood and investigating simulation-based inference such as developed by Ivanova et al., (2021); Kleinegesse and Gutmann, (2021); Kleinegesse et al., (2020). In addition, although in density-based BOED, we have shown that greedy approaches could outperform long-sighted reinforcement learning procedures, in a data-based setting, it would be interesting to investigate an extension to non myopic approaches such as Iqbal et al., (2024).
Acknowledgments
The authors thank the reviewers and area chair for their interesting and useful comments and the Inria Challenge project ROAD-AI for partial funding. This work was performed using HPC/AI resources from GENCI-IDRIS (Grant 2023-AD011014217R1). Pierre Alliez is supported by the French government, through the 3IA Côte d’Azur Investments in the Future project managed by the National Research Agency ANR-19-P3IA-0002.
@@ Appendix B Logarithmic pooling as a good importance sampling proposal
When considering importance sampling with a proposal distribution $q$ and a target distribution $p$, Chatterjee and Diaconis, (2018) proved that under certain conditions, the number of simulation draws required for both importance sampling and self normalized importance sampling (SNIS) estimators to have small L1 error with high probability was roughly $\exp(\text{KL}(p,q))$, see Theorem 1.2 in Chatterjee and Diaconis, (2018) for SNIS. Similarly, selecting a proposal distribution which minimizes the importance sampling estimator variance is equivalent to finding a distribution with small $\chi^{2}$-distance to $p$, see e.g. Appendix E of Minka, (2005). More generally, finding a good proposal $q$ is linked to the problem of minimizing $\alpha$-divergences or $f$-divergence between $p$ and $q$, which are jointly convex in $p$ and $q$, see Minka, (2005). In this work, we consider $\text{KL}(q,p)$ as a measure of proximity between $p$ and $q$. This choice is ultimately arbitrary but has the advantage of leading to an interpretable proposal with interesting sampling properties. To justify the pooled posterior $q_{{\boldsymbol{\xi}},N}$ in (9) and its use in (10), we then use Lemma 2 below to show that for $\sum_{i=1}^{N}\nu_{i}=1$, the distribution $q^{*}$ that minimizes the weighted sum of the KL against each posterior $p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}})$, i.e. $\sum_{i=1}^{N}\nu_{i}\text{KL}(q,p({\boldsymbol{\theta}}|{\mathb{y}}_{i},{\boldsymbol{\xi}}))$ is
@@
Example: as already mentioned our pooled posterior $q_{{\boldsymbol{\xi}},N}$ corresponds to the application of this result to $\rho=\sum_{i=1}^{N}\nu_{i}\delta_{{\mathb{y}}_{i}}$ with $\sum_{i=1}^{N}\nu_{i}=1$.
@@
Remark 3: The pooled posterior distribution can also be recovered as a constrained mean field solution. Indeed, it is easy to show that $q^{*}$ is also the measure that minimizes the KL between the joint distribution and a product form approximation where one of the factor is fixed to $\rho({\mathb{y}})$,
@@ C.1 Denoising diffusion models
where ${\boldsymbol{\epsilon}}\sim\mathcal{N}(\mathbf{0},{\mathb{I}})$ is a standard Gaussian random variable. Samples from $p_{0}$ are transformed to samples approximately from a standard Gaussian distribution after some large time $T$.
@@
where $t$ flows backwards from infinity to $t\!\!=\!\!0$ and $p_{t}$ is the distribution of $\tilde{{\boldsymbol{\theta}}}^{(t)}$ from (34). In practice, the process is started at some large finite $T$ assuming that ${\boldsymbol{\theta}}^{(T)}\sim{\cal N}(\mathbf{0},{\mathb{I}})$. In this Appendix, we rather consider increasing time $t$ from 0 to $T$, using that (37) can be equivalenty written as,
$d{\boldsymbol{\theta}}^{(t)}=\left[\frac{\beta(T-t)}{2}{\boldsymbol{\theta}}^{(t)}+\beta(T-t)\nabla_{\boldsymbol{\theta}}\log p_{T-t}({\boldsymbol{\theta}}^{(t)})\right]dt+\sqrt{\beta(T-t)}d{{\mathb{B}}}_{t}\;,$ (38)
which is now initialized with ${\boldsymbol{\theta}}^{(0)}\sim{\cal N}(\mathbf{0},{\mathb{I}})$. Solving this reverse SDE, the distribution of ${\boldsymbol{\theta}}^{(T)}$ is closed to $p_{0}$ for large $T$, which allows approximate sampling from $p_{0}$.
The score function $\nabla_{\boldsymbol{\theta}}\log p_{t}({\boldsymbol{\theta}})$ of the noisy data distribution at time $t$ is intractable and is then estimated by learning a neural network $s_{\phi}({\boldsymbol{\theta}},t)$ with parameters $\phi$. Score matching (Hyvärinen,, 2005) is a method to train $s_{\phi}$ by minimizing the following loss:
@@
${\boldsymbol{\theta}}^{(t+\Delta t)}={\boldsymbol{\theta}}^{(t)}+\frac{\beta(T-t)}{2}{\boldsymbol{\theta}}^{(t)}\Delta t+\beta(T-t)s_{\phi}({\boldsymbol{\theta}}^{(t)},T-t)\Delta t+\sqrt{\beta(T-t)\Delta t}{\boldsymbol{\epsilon}}\;.$ (41)
@@ C.2 Conditional Diffusion Models
$d{\boldsymbol{\theta}}^{(t)}=\left[\frac{\beta(T-t)}{2}{\boldsymbol{\theta}}^{(t)}+\beta(T-t)\nabla_{\boldsymbol{\theta}}\log p_{T-t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})\right]dt+\sqrt{\beta(T-t)}d{\mathb{B}}_{t},$ (43)
with the usual score $\nabla_{\boldsymbol{\theta}}\log p_{T-t}({\boldsymbol{\theta}}^{(t)})$ replaced by the conditionnal score $\nabla_{\boldsymbol{\theta}}\log p_{T-t}({\boldsymbol{\theta}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})$. The main objective of conditional SDE is to generate samples from the conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$ without retraining a new neural network $s_{\phi}$ for the new conditional score. Writing in terms of the forward process, the conditional score can be written using:
$\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}(\tilde{{\boldsymbol{\theta}}}^{(t)}|{\mathb{y}},{\boldsymbol{\xi}})=\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}({\mathb{y}}|\tilde{{\boldsymbol{\theta}}}^{(t)},{\boldsymbol{\xi}})+\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}(\tilde{{\boldsymbol{\theta}}}^{(t)})$ (44)
and we can leverage a pre-computed neural network $s_{\phi}$ that was trained to estimate the score $\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}(\tilde{{\boldsymbol{\theta}}}^{(t)})$ in the unconditional case. If we know how to evaluate the first term $\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}({\mathb{y}}|\tilde{{\boldsymbol{\theta}}}^{(t)},{\boldsymbol{\xi}})$, we can then run the reverse process (43) to generate samples from the conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$. Unfortunately, this term does not have a closed form expression. As a solution, Dou and Song, (2024) propose to approximate the intractable $\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}({\mathb{y}}|\tilde{{\boldsymbol{\theta}}}^{(t)},{\boldsymbol{\xi}})$ by the tractable $\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}({\mathb{y}}^{(t)}|\tilde{{\boldsymbol{\theta}}}^{(t)},{\boldsymbol{\xi}})$ where ${\mathb{y}}^{(t)}$ is a noisy version of ${\mathb{y}}$ at time $t$. Then, the following backward SDE can be run to generate samples from the conditional distribution $p({\boldsymbol{\theta}}|{\mathb{y}},{\boldsymbol{\xi}})$:
$d{\boldsymbol{\theta}}^{(t)}=$ $\left[\frac{\beta(T\!-\!t)}{2}{\boldsymbol{\theta}}^{(t)}+\beta(T\!-\!t)\nabla_{\theta}\log p_{T\!-\!t}({\mathb{y}}^{(T\!-\!t)}|{\boldsymbol{\theta}}^{(t)},{\boldsymbol{\xi}})+\beta(T\!-\!t)\nabla_{\theta}\log p_{T\!-\!t}({\boldsymbol{\theta}}^{(t)})\right]dt$
$+\sqrt{\beta(T\!-\!t)}d{\mathb{B}}_{t}$ (45)
@@
${\mathb{y}}^{(t)}={\mathb{A}}_{\boldsymbol{\xi}}\tilde{{\boldsymbol{\theta}}}^{(t)}+\sqrt{\bar{\alpha}_{t}}{\boldsymbol{\eta}}\;.$ (47)
We can then evaluate $\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}({\mathb{y}}^{(t)}|\tilde{{\boldsymbol{\theta}}}^{(t)},{\boldsymbol{\xi}})$ as :
$\nabla_{\tilde{{\boldsymbol{\theta}}}}\log p_{t}({\mathb{y}}^{(t)}|\tilde{{\boldsymbol{\theta}}}^{(t)},{\boldsymbol{\xi}})=\frac{1}{\sigma^{2}\bar{\alpha}_{t}}{\mathb{A}}_{\boldsymbol{\xi}}^{T}({\mathb{y}}^{(t)}-{\mathb{A}}_{\boldsymbol{\xi}}\tilde{{\boldsymbol{\theta}}}^{(t)})\;.$ (48)
C.3 Gradient estimation
When sampling is performed with a finite time horizon diffusion model, the first iterations of the corresponding sampling operator may provide too noisy samples which would result in gradient estimations with little information. One solution proposed by Marion et al., (2025) is to use a queuing trick which requires to store in memory a queue of samples. The memory burden is high and only acceptable for a limited number of particles. We propose a simpler solution, which consists in using Tweedie’s formula (Efron,, 2011) to perform a one-shot backward step, replacing a potentially too noisy ${\boldsymbol{\theta}}^{(t)}$ by the conditional mean $\mathbb{E}[\tilde{{\boldsymbol{\theta}}}^{(0)}|\tilde{{\boldsymbol{\theta}}}^{(T-t)}={\boldsymbol{\theta}}^{(t)}]$, which can be interpreted as its prediction at time 0. The Tweedie’s formula provides this prediction in close-form,
$\hat{{\boldsymbol{\theta}}}^{(t)}$ $=\mathbb{E}[\tilde{{\boldsymbol{\theta}}}^{(0)}|\tilde{{\boldsymbol{\theta}}}^{(T-t)}={\boldsymbol{\theta}}^{(t)}]=\frac{{\boldsymbol{\theta}}^{(t)}+(1-\bar{\alpha}_{T-t})\nabla_{\boldsymbol{\theta}}\log p_{T-t}({\boldsymbol{\theta}}^{(t)})}{\sqrt{\bar{\alpha}_{T-t}}}\;.$ (49)
Gradients are then computed using the $\hat{{\boldsymbol{\theta}}}^{(t)}$’s values, while ${\boldsymbol{\theta}}^{(t)}$ is updated into ${\boldsymbol{\theta}}^{(t+1)}$ from the backward SDE, as mentioned above. This is only really impactful for small $t$ as for large $t$, $\hat{{\boldsymbol{\theta}}}^{(t)}$ and ${\boldsymbol{\theta}}^{(t)}$ get closer. This extra computation does not add cost as (49) uses a score value that is already computed for (C.2).